# Using nonassociative algebras to classify skew polycyclic codes up to isometry and equivalence

### v1

@@ 5.2. Skew polycyclic codes over finite fields
(i) Let $n\nmid m$. Then the number of different Chen equivalence classes of skew $(f,\sigma)$-polycyclic codes of length $m$ is at most
$N=\gcd([m]_{s},p^{r}-1).$
(ii) Let $n|m$. Then the number of different Chen equivalence classes of skew $(f,\sigma)$-polycyclic codes of length $m$ is at most
$N=\left(1-\frac{1}{[n]_{s}}\right)\gcd([m]_{s},p^{r}-1)+\gcd([m]_{s},p^{r}-1)/[n]_{s}.$
In this count, $\left(1-\frac{1}{[n]_{s}}\right)\gcd([m]_{s},p^{r}-1)$ equivalence classes belong to skew polynomials $f$ that have constant coefficient $a_{0}\notin\mathbb{F}_{p^{s}}$.
Proof.
$(i)$ If $n\nmid m$, then the algebras $\mathbb{F}_{p^{r}}[t;\sigma]/\mathbb{F}_{p^{r}}[t;\sigma](t^{m}-a)$ are not associative and there are $\gcd([m]_{s},p^{r}-1)$ classes of Chen equivalent $(\sigma,a)$-constacyclic codes. This immediately yields the assertion.
$(ii)$ If $n|m$, then $N_{m}^{\sigma}(\mathbb{F}_{p^{r}})\subset\mathbb{F}_{p^{s}}$, so that every coset of $N_{m}^{\sigma}(\mathbb{F}_{p^{r}})$ is either entirely contained in $\mathbb{F}_{p^{s}}$ or entirely disjoint from it (so corresponds to a nonassociative Petit algebra). We know that $\mathbb{F}_{p^{r}}[t;\sigma]/\mathbb{F}_{p^{r}}[t;\sigma](t^{m}-a)$ is a proper nonassociative algebra if and only if $a\in\mathbb{F}_{p^{r}}\setminus\mathbb{F}_{p^{s}}$. Turning to the proof of [NevPu2025, Theorem 5.6], the total number of cosets contained in $\mathbb{F}_{p^{s}}$, and thus corresponding to associative algebras where $a\in\mathbb{F}_{p^{s}}$, is equal to $\frac{p^{s}-1}{(p^{r}-1)/w}=\frac{w}{[n]_{s}}$. Subtracting these $\frac{w}{[n]_{s}}$ cosets from the total number of cosets gives the number of Chen equivalence classes for nonassociative algebra that is $N=w-w/[n]_{s}$. This yields the assertions. ∎
5.3. When two classes cannot be equivalent
If we want to know when skew polycyclic codes cannot be equivalent, there are some easy conditions we can check first. We are looking at classes ${\bf C}_{f}$ and ${\bf C}_{h}$ of skew polycyclic codes of length $m$ over $S$.
@@ 5.4. Conclusions and further work
We initiated the study of skew polycyclic codes up to equivalence, identified skew polycyclic codes which are the same in terms of protecting the underlying information, and obtained some catalogues of codes up to isometries and equivalences with tools for de-duplicating them, identifying gaps on which further research may be focused. Our equivalence relations yield tighter classifications of isometric and equivalent codes that previously used ones.
Our next goal is to now look at special $S$, like finite fields or rings, and explicitly compute the equivalence classes. The isometry classes for skew polycyclic codes seem to be hard to find, however, the special case of skew constacyclic codes seems tractable.
@@ Acknowledgments
This paper was written during the second author’s stay as a Simons Professor in Residence at the University of Ottawa. She gratefully acknowledges the support of CRM and the Simons Foundation.

### v2

@@ 6. Skew polycyclic codes over finite fields
Let $n\geq m-1$ and assume that $f$ does not generate a two-sided ideal in $\mathbb{F}_{q}[t;\sigma]$. The exact number of Chen isometry classes for codes of length $m$ over $\mathbb{F}_{q}$ computed in [OaHA25, Theorem 4.20] is
$N=\frac{(q-1)^{m_{0}}}{lcm_{j\in\{i_{0},i_{1},\dots,i_{m_{0}}-1\}}\big(\frac{q-1}{gcd(|m-j|_{s},q-1)}\big)}$
where the $j$ run over all the $m_{0}$ different nonzero coefficients of $f$.
7. When two classes cannot be equivalent
If we want to have some easy to check criteria for when skew polycyclic codes cannot be equivalent, there are some easy conditions we can check first. We are looking at classes ${\bf C}_{f}$ and ${\bf C}_{h}$ of skew polycyclic codes of length $m$ over $S$.
@@ 7.1. Conclusions and further work
We initiated the study of skew polycyclic codes up to equivalence, identified skew polycyclic codes which are the same in terms of protecting the underlying information, and obtained some catalogues of skew constacyclic codes up to isometries and equivalences with tools for de-duplicating them, identifying gaps on which further research may be focused. Our equivalence relations yield tighter classifications of isometric and equivalent codes that previously used ones.
Our next goal is to now look at special rings $S$, like finite fields or rings, and explicitly compute the equivalence classes. The isometry classes for skew polycyclic codes seem to be hard to find, however, the special case of skew constacyclic codes seems tractable.
@@ Acknowledgments
This paper was written during the second author’s stay as a Simons Professor in Residence at the University of Ottawa. She gratefully acknowledges the support of CRM and the Simons Foundation, and thanks the Department of Mathematics and Statistics for its hospitality and its congenial and inspiring atmosphere, especially M Nevins for lots of inspriring conversations.